# Deep Reinforcement Learning for Real-Time Drone Routing in Post-Disaster Road Assessment Without Domain Knowledge

### v1

@@
[…]
Deep Reinforcement Learning for Real-Time Drone Routing in Post-Disaster Road Assessment Without Domain Knowledge
[…]
Rapid post-disaster road damage assessment is critical for effective emergency response, yet traditional optimization methods suffer from excessive computational time and require domain knowledge for algorithm design, making them unsuitable for time-sensitive disaster scenarios. This study proposes an attention-based encoder-decoder model (AEDM) for real-time drone routing decision in post-disaster road damage assessment. The method employs deep reinforcement learning to determine high-quality drone assessment routes without requiring algorithmic design knowledge. A network transformation method is developed to convert link-based routing problems into equivalent node-based formulations, while a synthetic road network generation technique addresses the scarcity of large-scale training datasets. The model is trained using policy optimization with multiple optima (POMO) with multi-task learning capabilities to handle diverse parameter combinations. Experimental results demonstrate two key strengths of AEDM: it outperforms commercial solvers by 16–69% in solution quality and achieves real-time inference (1–2 seconds) versus 100–2,000 seconds for traditional methods. The model exhibits strong generalization across varying problem scales, drone numbers, and time constraints, consistently outperforming baseline methods on unseen parameter distributions and real-world road networks. The proposed method effectively balances computational efficiency with solution quality, making it particularly suitable for time-critical disaster response applications where rapid decision-making is essential for saving lives.
[…]
Road damage assessment, Drone routing decision, Deep reinforcement learning, Attention-based encoder-decoder, Network transformation
@@ 1 Introduction
In recent years, the frequency and intensity of both natural disasters (e.g., earthquakes, hurricanes, and wildfires) and man-made crises (e.g., conflicts) have increased significantly, posing severe threats to human lives, property, and economic stability (Avishan et al. 2023). For example, the 2011 Japan earthquake resulted in over 15,000 deaths and approximately \$235 billion in economic losses, while the 2023 Turkey-Syria earthquake caused more than 50,000 fatalities (Sun et al. 2022). Between 2016 and 2020, disasters affected over 602 million people globally (Yang et al. 2023). Timely post-disaster emergency responses that include the delivery of essential supplies such as food, water, and medical aid are critical for saving lives (Organization et al. 2000). However, these efforts are often hindered by uncertainties in transportation infrastructure disruptions immediately following a disaster. Such uncertainties encompass the extent of road network damage, current operability, and feasibility of short-term restoration. Thus, rapid assessment of road network damage is essential to mitigate this uncertainty and enable effective humanitarian interventions.
@@
Unmanned aerial vehicles, or drones, have emerged as a transformative tool for post-disaster damage assessment, offering distinct advantages over traditional methods (Enayati et al. 2023, Shi et al. 2024, Steenbergen et al. 2025). Their ability to capture high-resolution imagery at lower operational costs than satellites or helicopters makes them well-suited for rapid deployment in disaster zones (Zhang et al. 2023). Equipped with advanced sensors such as high-definition cameras, drones can systematically collect detailed data on road network conditions, including damage type and severity (e.g., cracks, landslides or complete collapses), current trafficability, and feasibility of short-term repairs. For instance, during the 2013 Typhoon Haiyan response in the Philippines, drones rapidly mapped affected areas and identified blocked roads, enabling relief teams to prioritize clearance efforts (Turk 2014). In Nepal’s 2015 earthquake recovery, thermal drones assisted in locating survivors and assessing hard-to-reach mountain roads, showcasing their versatility in complex terrains (Abuali and Ahmed 2025). In post-disaster scenarios, traditional transportation options like cars, motorcycles, and even pedestrian access often become impassable and hazardous (Zhang et al. 2023). Roads may be collapsed or blocked by landslides, making ground-based access extremely difficult or even life-threatening. Drones, however, can operate freely in such hazardous environments, bypassing damaged infrastructure to collect critical information (Enayati et al. 2023, Shi et al. 2024, Steenbergen et al. 2025). This agility is pivotal in disaster response, where time sensitivity means timely information can determine survival outcomes. Building on these validated applications, this study proposes deploying a fleet of drones for rapid post-disaster road damage assessment.
@@
Post-disaster road damage assessment via drone encompasses two critical timeframes: 1) the duration required to determine the routes for road damage assessment, and 2) the time taken by each drone to traverse its designated route. The first timeframe focuses on rapidly determining drone routing paths to enable efficient drone deployment. Existing literature tends to rely on traditional methodologies, including heuristics (e.g., genetic algorithms and simulated annealing) or exact methods (e.g., branch-and-price and column generation), to solve this routing problem (Huang et al. 2013, Balcik and Yanıkoğlu 2020). However, these methods suffer from three key limitations. First, their computational time increases significantly with problem size, making them ill-suited for time-critical disaster scenarios where delays in providing solutions can be unacceptable (Luo et al. 2023, Liu et al. 2024b). Second, developing a universally efficient algorithm to cope with such a drone routing problem is challenging. Such methods typically require extensive domain-specific expertise to engineer problem-tailored heuristics or exact formulations, which leads to prohibitive development costs and prolonged deployment cycles. Third, the problem’s constraints evolve rapidly with changing environments (Kwon et al. 2020, Liu et al. 2024b), thus necessitating frequent redesign of specialized algorithms.
@@
Advancements in computer vision (CV) and large language models (LLMs) have revolutionized their respective fields: expert-driven manual feature engineering has been supplanted by automated end-to-end deep learning, enabling models to learn complex patterns and generalize across diverse scenarios without heavy human intervention (Brown et al. 2020, Radford et al. 2021, Luo et al. 2023, Zhou et al. 2024, Liu et al. 2024a, Berto et al. 2024, Huang et al. 2025). Drawing inspiration from this paradigm shift, this study adopts a similar philosophy by proposing a deep reinforcement learning (DRL) model. Like CV and LLM that autonomously extract meaningful representations from raw data, the DRL model eliminates the need for domain-specific algorithm design by learning optimal routing strategies directly from problem instances. This allows it to determine high-quality drone assessment routes in real time, adapting to varying disaster scenarios and outperforming traditional algorithms by a significant margin. The second timeframe ensures timely completion of assessments from route initiation to drone return. The proposed model explicitly incorporates this timeliness constraint, optimizing both route planning and completion time within a predefined limit.
To achieve the above objectives, this study presents three contributions:
  • We propose an attention-based encoder-decoder model (AEDM) for real-time decision-making in drone assessment route planning. Unlike traditional methods that rely on domain knowledge for algorithm design, AEDM operates autonomously. Experimental results demonstrate that it outperforms commercial solvers by a significant margin, with improvements ranging from 16–69%. More crucially, AEDM maintains real-time solving capability for large-scale problems. In contrast, commercial solvers fail to yield solutions within a reasonable time frame for such complex scenarios, highlighting the model’s superiority in time-critical post-disaster assessment tasks.
  • We propose a simple yet efficient network transformation approach. This method integrates two distinct networks into a unified structure, enabling the conversion of a link-based routing problem into an equivalent node-based routing problem.
  • We propose a synthetic road network generation method tailored for DRL in road network-related problems. This method addresses the scarcity of real-world large-scale road network datasets suitable for training advanced learning models.
@@ 2 Related work
In prior road damage assessment practices, four primary approaches have been employed. The first approach entails field surveys, which consist of on-site investigative activities such as community mapping and transect walks (Maya 2013). While these methods can provide detailed data, they are time-consuming, labor-intensive, and often fail to deliver timely insights during the critical early phases of disaster response (Tatham 2009). The second approach relies on remote sensing and geographic information system-based methods. These utilize satellite imagery to rapidly identify impacted areas. However, satellite revisit intervals (ranging from days to weeks) and resolution constraints limit their capacity to support prompt responses and detect fine-grained damages (Bravo et al. 2019). The third approach makes use of traffic data and sensor networks. It leverages global positioning systems, cameras, and embedded sensors to infer real-time road conditions. Yet, the limited deployment of such sensors in remote or underdeveloped regions undermines their feasibility for large-scale disaster scenarios (Yang et al. 2023). The fourth and final approach features drones. Drones emerge as a highly promising solution, owing to their high mobility, rapid deployability, and ability to capture high-resolution data in hazardous environments (Agatz et al. 2018, Zhang et al. 2023).
@@
Recent literature has explored drone applications in post-disaster assessment. For example, Huang et al. (2013) formulated an assessment routing problem focusing on minimizing the total response time for visiting affected communities. Balcik (2017) and Balcik and Yanıkoğlu (2020) extended this to optimize site selection and routing under uncertain travel conditions, using traditional optimization techniques such as mixed-integer linear programming (MILP) solvers, genetic algorithms, or tabu search. However, these methods require significant computational time as problem size increases, making real-time deployment challenging. Oruc and Kara (2018) combined drones and ground vehicles to optimize the assessment process but constrained drone movement to the physical road network, limiting flexibility. Zhang et al. (2023) introduced a robust team-orienteering model solved via a branch-and-price approach, which remains computationally intensive for larger problems. Meanwhile, Adsanver et al. (2025) applied a multistage framework for grid damage assessment, but their heuristic-based models exhibit increased solution time with problem complexity. Table 1 presents a comparative overview of these methods, highlighting key differences in solution approaches, computational performance, and reliance on domain knowledge. It shows that traditional methods, whether heuristic (e.g., genetic algorithms and tabu search) or exact (e.g., branch-and-price), exhibit solution time that increases significantly with problem size even for moderate scales (e.g., 100+ nodes) and inherently require extensive domain-specific expertise for algorithm design. In contrast, our proposed DRL-based approach achieves real-time performance while handling larger networks (up to 1,000 nodes) and eliminates the need for domain knowledge, addressing the critical time sensitivity and expertise barriers of disaster response.
@@
Given the time sensitivity of disaster response, the determination of assessment routes for each drone post-disaster must be as rapid as possible, ideally in real time. Heuristic algorithms, which aim to balance computational efficiency and solution quality, exhibit three notable limitations: First, they require substantial domain expertise to design effective rules, rendering the process time-consuming and costly (Liu et al. 2022, Luo et al. 2023, Zhou et al. 2024, Liu et al. 2024a). Second, they lack flexibility when faced with environmental changes or varying problem types, necessitating frequent adjustments (Zhang et al. 2020, Berto et al. 2024). Third, while performing adequately on smaller-scale problems, they incur longer solution time for larger scenarios, reducing computational efficiency and making them ill-suited for real-time applications (Chen et al. 2022, Arishi and Krishnan 2023). As a complement to these existing approaches, this study specifically addresses the urgency of post-disaster assessment by proposing a DRL model that determines high-quality drone routing plans in real time without domain knowledge. Our approach leverages advances in artificial intelligence (AI) to overcome the scalability and flexibility limitations inherent in classical optimization and heuristic methods, thereby enabling real-time and efficient damage assessment in complex post-disaster environments.
@@
Authors Assessment tools Objectives Solutions Method Case study size Computational time Domain knowledge
Huang et al. (2013) Routes teams Minimize the sum of arrival time at all nodes Continuous approximation approach Heuristic 10 and 20 teams Increase significantly with the case size Required
Balcik (2017) Routes teams Cover nodes with different characteristics in a balanced way Tabu search Heuristic 94 nodes Increase significantly with the case size Required
Balcik and Yanıkoğlu (2020) Routes teams Cover nodes with different characteristics in a balanced way under uncertain travel time Tabu search Heuristic 94 nodes Increase significantly with the case size Required
Oruc and Kara (2018) Drones and motorcycles Maximize the total value added by the assessment of road links and maximize the total profit generated by assessing points of interest (nodes) Base route heuristic Heuristic 44 nodes Increase significantly with the case size Required
Glock and Meyer (2020) Drones Maximize the total informativeness of sampled locations within drones flight time limits Two-phase multistart adaptive large neighborhood search Heuristic Up to 625 nodes Increase significantly with the case size Required
Zhang et al. (2023) Drones Maximize the collected link informative reward within a predefined time limit Branch-and-price algorithm Exact Up to 107 nodes Increase significantly with the case size Required
Yin et al. (2023) Trucks and drones Minimize sum of dispatching cost, transport cost, and worst penalty cost of tardiness under demand and travel time uncertainties Branch-and-price-and-cut algorithm Exact Up to 45 nodes Increase significantly with the case size Required
Morandi et al. (2024) Truck and drones Maximize total collected prize from nodes (visited by truck or served by drones) within a predefined time limit Branch-and-cut algorithm Exact Up to 50 nodes Increase significantly with the case size Required
Adsanver et al. (2025) Drones Maximize the total priority scores of the visited grids and penalizes deviations from the maximum coverage target Variable neighborhood descent Heuristic Up to 148 nodes Increase significantly with the case size Required
This study Drones Maximize the collected link informative reward within a predefined time limit DRL Heuristic Up to 1,000 nodes Real-time Not required
@@ 3 Problem definition
In the immediate aftermath of a disaster, the highly uncertain extent of road network damage underscores the critical need to rapidly assess network usability before executing emergency supply deployments and evacuation efforts. This problem is formally framed as follows. Given a road network modeled as a graph $G=(N,A)$ (where $N$ denotes the set of road nodes, encompassing intersections in urban settings or intersections/villages in rural contexts, and $A$ represents the links between these nodes) and a fleet of drones, the objective is to determine the assessment routes for each drone. These routes aim to maximize the total value of damage-related information collected across the network links. Drones, outfitted with high-definition cameras, are deployed to evaluate the damage status of each link. Due to a drone’s operating altitude of 120 meters and its visual range (forming a circular area with a radius of approximately 75–80 meters), simultaneously inspecting multiple roads is impractical even within urban environments (Zhang et al. 2023). Consequently, a bidirectional road link can only be assessed if a drone flies directly along it.
@@
Figure 1: Dual Network: Road Links (Assessment) vs. Direct Routes (Transit)

### v2

@@
Abstract
Rapid post-disaster road damage assessment is critical for effective emergency response, yet traditional optimization methods suffer from excessive computational time and require domain knowledge for algorithm design, making them unsuitable for time-sensitive disaster scenarios. This study proposes an attention-based encoder-decoder model (AEDM) for rapid drone routing decision in post-disaster road damage assessment. The method employs deep reinforcement learning to determine high-quality drone assessment routes without requiring algorithmic design knowledge. A network transformation method is developed to convert link-based routing problems into equivalent node-based formulations, while a synthetic road network generation technique addresses the scarcity of large-scale training datasets. The model is trained using policy optimization with multiple optima (POMO) with multi-task learning capabilities to handle diverse parameter combinations. Experimental results demonstrate two key strengths of AEDM: it outperforms commercial solvers by 20–71% and traditional heuristics by 23–35% in solution quality, while achieving rapid inference (1–2 seconds) versus 100–2,000 seconds for traditional methods. The model exhibits strong generalization across varying problem scales, drone numbers, and time constraints, consistently outperforming baseline methods on unseen parameter distributions and real-world road networks. The proposed method effectively balances computational efficiency with solution quality, making it particularly suitable for time-critical disaster response applications where rapid decision-making is essential for saving lives. The source code for AEDM is publicly available at https://github.com/PJ-HTU/AEDM-for-Post-disaster-road-assessment.
keywords:
@@ 1 Introduction
In recent years, the frequency and intensity of both natural disasters (e.g., earthquakes, hurricanes, and wildfires) and man-made crises (e.g., conflicts) have increased significantly, posing severe threats to human lives, property, and economic stability (avishan2023humanitarian). For example, the 2011 Japan earthquake resulted in over 15,000 deaths and approximately \$235 billion in economic losses, while the 2023 Turkey-Syria earthquake caused more than 50,000 fatalities (sun2022novel). Between 2016 and 2020, disasters affected over 602 million people globally (yang2023distributionally). Timely post-disaster emergency responses that include the delivery of essential supplies such as food, water, and medical aid are critical for saving lives (world2000natural). However, these efforts are often hindered by uncertainties in transportation infrastructure disruptions immediately following a disaster. Such uncertainties encompass the extent of road network damage, current operability, and feasibility of short-term restoration. Thus, rapid assessment of road network damage is essential to mitigate this uncertainty and enable effective humanitarian interventions.
@@
Unmanned aerial vehicles, or drones, have emerged as a transformative tool for post-disaster damage assessment, offering distinct advantages over traditional methods (enayati2023multimodal; shi2024optimal; van2025stochastic). Their ability to capture high-resolution imagery at lower operational costs than satellites or helicopters makes them well-suited for rapid deployment in disaster zones (zhang2023robust). Equipped with advanced sensors such as high-definition cameras, drones can systematically collect detailed data on road network conditions, including damage type and severity (e.g., cracks, landslides or complete collapses), current trafficability, and feasibility of short-term repairs. For instance, during the 2013 Typhoon Haiyan response in the Philippines, drones rapidly mapped affected areas and identified blocked roads, enabling relief teams to prioritize clearance efforts (turk2014drones). In Nepal’s 2015 earthquake recovery, thermal drones assisted in locating survivors and assessing hard-to-reach mountain roads, showcasing their versatility in complex terrains (abuali2025innovative). In post-disaster scenarios, traditional transportation options like cars, motorcycles, and even pedestrian access often become impassable and hazardous (zhang2023robust). Roads may be collapsed or blocked by landslides, making ground-based access extremely difficult or even life-threatening. Drones, however, can operate freely in such hazardous environments, bypassing damaged infrastructure to collect critical information (enayati2023multimodal; shi2024optimal; van2025stochastic). This agility is pivotal in disaster response, where time sensitivity means timely information can determine survival outcomes. Building on these validated applications, this study proposes deploying a fleet of drones for rapid post-disaster road damage assessment.
@@
Post-disaster road damage assessment via drone encompasses two critical timeframes: 1) the duration required to determine the routes for road damage assessment, and 2) the time taken by each drone to traverse its designated route. The first timeframe focuses on rapidly determining drone routing paths to enable efficient drone deployment. Existing literature tends to rely on traditional methodologies, including heuristics (e.g., genetic algorithms and simulated annealing) or exact methods (e.g., branch-and-price and column generation), to solve this routing problem (huang2013continuous; balcik2020robust). However, these methods suffer from three key limitations. First, their computational time increases significantly with problem size, making them ill-suited for time-critical disaster scenarios where delays in providing solutions can be unacceptable (luo2023neural; liu2024evolution). Second, developing a universally efficient algorithm to cope with such a drone routing problem is challenging. Such methods typically require extensive domain-specific expertise to engineer problem-tailored heuristics or exact formulations, which leads to prohibitive development costs and prolonged deployment cycles. Third, the problem’s constraints evolve rapidly with changing environments (kwon2020pomo; liu2024evolution), thus necessitating frequent redesign of specialized algorithms.
@@
Advancements in computer vision (CV) and large language models (LLMs) have revolutionized their respective fields: expert-driven manual feature engineering has been supplanted by automated end-to-end deep learning, enabling models to learn complex patterns and generalize across diverse scenarios without heavy human intervention (brown2020language; radford2021learning; luo2023neural; zhou2024mvmoe; liu2024multi; berto2024routefinder; huang2025orlm). Drawing inspiration from this paradigm shift, this study adopts a similar philosophy by proposing a deep reinforcement learning (DRL) model. Like CV and LLM that autonomously extract meaningful representations from raw data, the DRL model eliminates the need for domain-specific algorithm design by learning optimal routing strategies directly from problem instances. This allows it to determine high-quality drone assessment routes within seconds, adapting to varying disaster scenarios and outperforming traditional algorithms by a significant margin. The second timeframe ensures timely completion of assessments from route initiation to drone return. The proposed model explicitly incorporates this timeliness constraint, optimizing both route planning and completion time within a predefined limit. To achieve the above objectives, this study presents the following contributions:
  • We propose an attention-based encoder-decoder model (AEDM) for rapid drone route planning in post-disaster road assessment. Unlike traditional methods requiring domain-specific algorithm design, AEDM learns optimal routing strategies autonomously through deep reinforcement learning.
  • We develop a network transformation and synthetic data generation framework that converts link-based routing into node-based formulation while generating diverse training instances, addressing the dual challenges of problem complexity and dataset scarcity.
  • We propose a multi-task reinforcement learning framework with exponential moving average based reward normalization that enables simultaneous training across varying drone numbers and time constraints, eliminating the need for separate models per parameter combination.
  • We conduct extensive experiments demonstrating that AEDM achieves 20–71% improvements over commercial solvers and 23–35% over traditional heuristics while maintaining rapid inference (1–2 seconds), with strong generalization to unseen parameters and real-world networks.
@@ 2 Related work
In prior road damage assessment practices, four primary approaches have been employed. The first approach entails field surveys, which consist of on-site investigative activities such as community mapping and transect walks (maya2013rapid). While these methods can provide detailed data, they are time-consuming, labor-intensive, and often fail to deliver timely insights during the critical early phases of disaster response (tatham2009investigation). The second approach relies on remote sensing and geographic information system-based methods. These utilize satellite imagery to rapidly identify impacted areas. However, satellite revisit intervals (ranging from days to weeks) and resolution constraints limit their capacity to support prompt responses and detect fine-grained damages (bravo2019use). The third approach makes use of traffic data and sensor networks. It leverages global positioning systems, cameras, and embedded sensors to infer real-time road conditions. Yet, the limited deployment of such sensors in remote or underdeveloped regions undermines their feasibility for large-scale disaster scenarios (yang2023distributionally). The fourth and final approach features drones. Drones emerge as a highly promising solution, owing to their high mobility, rapid deployability, and ability to capture high-resolution data in hazardous environments (agatz2018optimization; zhang2023robust).
@@
Recent literature has explored drone applications in post-disaster assessment. For example, huang2013continuous formulated an assessment routing problem focusing on minimizing the total response time for visiting affected communities. balcik2017site and balcik2020robust extended this to optimize site selection and routing under uncertain travel conditions, using traditional optimization techniques such as mixed-integer linear programming (MILP) solvers, genetic algorithms, or tabu search. However, these methods require significant computational time as problem size increases, making real-time deployment challenging. oruc2018post combined drones and ground vehicles to optimize the assessment process but constrained drone movement to the physical road network, limiting flexibility. zhang2023robust introduced a robust team-orienteering model solved via a branch-and-price approach, which remains computationally intensive for larger problems. Meanwhile, adsanver2025predictive applied a multistage framework for grid damage assessment, but their heuristic-based models exhibit increased solution time with problem complexity. Table 1 presents a comparative overview of these methods, highlighting key differences in solution approaches, computational performance, and reliance on domain knowledge. It shows that traditional methods, whether heuristic (e.g., genetic algorithms and tabu search) or exact (e.g., branch-and-price), exhibit solution time that increases significantly with problem size even for moderate scales (e.g., 100+ nodes) and inherently require extensive domain-specific expertise for algorithm design. In contrast, our proposed DRL-based approach achieves rapid performance while handling larger networks (up to 1,000 nodes) and eliminates the need for domain knowledge, addressing the critical time sensitivity and expertise barriers of disaster response.
@@
Given the time sensitivity of disaster response, the determination of assessment routes for each drone post-disaster must be as rapid as possible, ideally in real time. Heuristic algorithms, which aim to balance computational efficiency and solution quality, exhibit three notable limitations: First, they require substantial domain expertise to design effective rules, rendering the process time-consuming and costly (liu2022flying; luo2023neural; zhou2024mvmoe; liu2024multi). Second, they lack flexibility when faced with environmental changes or varying problem types, necessitating frequent adjustments (zhang2020multi; berto2024routefinder). Third, while performing adequately on smaller-scale problems, they incur longer solution time for larger scenarios, reducing computational efficiency and making them ill-suited for rapid applications (chen2022deep; arishi2023multi). As a complement to these existing approaches, this study specifically addresses the urgency of post-disaster assessment by proposing a DRL model that determines high-quality drone routing plans within seconds without domain knowledge. Our approach leverages advances in artificial intelligence (AI) to overcome the scalability and flexibility limitations inherent in classical optimization and heuristic methods, thereby enabling rapid and efficient damage assessment in complex post-disaster environments.
@@
Authors Tools Objectives Solution Type Scale Time Expertise
huang2013continuous Route teams Minimize arrival time Continuous approx. Heuristic 10–20 teams Increases Required
balcik2017site Route teams Balanced coverage Tabu search Heuristic 94 nodes Increases Required
balcik2020robust Route teams Balanced coverage (uncertain) Tabu search Heuristic 94 nodes Increases Required
oruc2018post Drones & bikes Max. link & node value Base route Heuristic 44 nodes Increases Required
glock2020mission Drones Max. informativeness ALNSa Heuristic 625 nodes Increases Required
zhang2023robust Drones Max. info. within time Branch-&-price Exact 107 nodes Increases Required
yin2023robust Trucks & drones Min. costs (uncertain) Branch-price-cut Exact 45 nodes Increases Required
morandi2024orienteering Truck & drones Max. prize within time Branch-&-cut Exact 50 nodes Increases Required
adsanver2025predictive Drones Max. priority scores VNDb Heuristic 148 nodes Increases Required
This study Drones Max. info. within time DRL (AEDM) Learning 1,000 nodes 1–2s Not req.
Note: aALNS = Adaptive Large Neighborhood Search; bVND = Variable Neighborhood Descent. “Increases” indicates computation time increases significantly with problem size. Max. = Maximize; Min. = Minimize.
@@ 3 Problem definition
In the immediate aftermath of a disaster, the highly uncertain extent of road network damage underscores the critical need to rapidly assess network usability before executing emergency supply deployments and evacuation efforts. This problem is formally framed as follows. Given a road network modeled as a graph $G=(N,A)$ (where $N$ denotes the set of road nodes, encompassing intersections in urban settings or intersections/villages in rural contexts, and $A$ represents the links between these nodes) and a fleet of drones, the objective is to determine the assessment routes for each drone. These routes aim to maximize the total value of damage-related information collected across the network links. Drones, outfitted with high-definition cameras, are deployed to evaluate the damage status of each link. Due to a drone’s operating altitude of 120 meters and its visual range (forming a circular area with a radius of approximately 75–80 meters), simultaneously inspecting multiple roads is impractical even within urban environments (zhang2023robust). Consequently, a bidirectional road link can only be assessed if a drone flies directly along it.
@@
Figure 1: Dual Network with Road Links for Assessment and Direct Routes for Transit